OpenStudy (anonymous):

You draw two marbles from a bag containing 3 red marbles and 6 blue marbles what is the probability that both marbles are red

OpenStudy (anonymous):

Probability = favorable cases / total cases favorable cases = choose 2 balls out three red balls = \[\left(\begin{matrix}3 \\ 2\end{matrix}\right) = 3\] Total cases = choose 2 balls out of all 6 balls = \[\left(\begin{matrix}6 \\ 2\end{matrix}\right) = 15\] Probability = 3/15 = 1/5

OpenStudy (vincent-lyon.fr):

Total cases should be 2 balls out of 9 balls: \(\left(\begin{matrix}9 \\ 2\end{matrix}\right) = 36\)

OpenStudy (anonymous):

@Vincent-Lyon.Fr Oh yes! Total cases should be 36.Thanks So the probability is 1/12. .